OpenStudy (anonymous):

given the linear equation y=-3/2x+1 find the y-coordinatesof the points (-4, )(-2, ) and(6, ).

OpenStudy (anonymous):

\[y=\frac{-3}{2}(-4)+1\] \[y=\frac{-3}{2}(-2)+1\] \[y=\frac{-3}{2}*6+1\]

OpenStudy (anonymous):

so my answers are( -4,-3),(-2,-1),&(6,7)

OpenStudy (anonymous):

Order of operations. You have to multiply -3/2 * x then add the one

OpenStudy (anonymous):

i dont have proper calculator :(

OpenStudy (anonymous):

\[\frac{-3}{2}*\frac{-4}{1}\]

OpenStudy (anonymous):

Simple multiplications. Shouldn't need a calculator to get these ones.

OpenStudy (anonymous):

-3 * -4 over 2 * 1

OpenStudy (anonymous):

15/2

OpenStudy (anonymous):

close 3 times 4 is 12 not 15 so 12/2 which is 6 because 2 goes into 12 6 times.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

y=6+1 so your first coordinate it (-4, 7)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the next one is 4 ?

OpenStudy (anonymous):

\[y=\frac{-3}{2}*\frac{-2}{1}+1\]

OpenStudy (anonymous):

Yes :]

OpenStudy (anonymous):

and then 10

OpenStudy (anonymous):

\[y=\frac{-3}{2}*\frac{6}{1}+1\]

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

oh a negative

OpenStudy (anonymous):

not -10 though

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

-3 * 6 over 2 * 1 -18/2 2 goes into -18 negative nine times. -9 + 1 = ?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

yep, thats all of them

OpenStudy (anonymous):

Thanks so much!=)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

U help is really appriciated